OpenStudy (anonymous):

How do you solve for y: 3x – 4y = 8

OpenStudy (tyteen4a03):

Throw everything that is not y to the right side.

OpenStudy (anonymous):

@tyteen4a03 lol good one

OpenStudy (anonymous):

3x - 4y = 8 -3x -3x -4y = -3x + 8 divide both sides by -4 y = 3/4x -2

OpenStudy (anonymous):

@Luis_Rivera jajaja. the only serious question I saw today was at 6am (the problem is still incomplete). all others are lots of fun

OpenStudy (anonymous):

:)